# Optimal Low-Depth Quantum Signal-Processing Phase Estimation

### v1

@@ Results
The key behind the success of QSPE is the isolation of $\theta$ and $\varphi$ estimations in Fourier space. This enables the robustness of individual angle estimation against the error and noise in another angle. A prior art that is widely used in the gate calibration in Google’s superconducting platform is periodic calibration or Floquet calibration [2, 23]. Periodic calibration measures the transition probability between tensor product states |$\ket{10}$ and |$\ket{01}$ of a periodic quantum circuit with $d$ $U$-gates inside. This differs from our QSPE method which initializes Bell states, though the main body of the quantum circuit is the same. To provide an estimation of angles, periodic calibration uses a black-box optimization to minimize the distance between the parametric ansatz and experimentally measured values. Periodic calibration is based on RPE [14, 23, 2] and generalizes RPE to the estimation of multiple angles. Though RPE provably saturates the Heisenberg limit, the actual performance of periodic calibration highly relies on the satisfaction of the so-called phase-matching condition, namely, $\omega=\varphi$. Previous experiments suggest that the violation of the condition would lower the estimation accuracy of $\theta$ angle by a few magnitudes. Because the phase angle $\varphi$ is vulnerable to time-dependent drift errors, the uncertainty of $\varphi$ ultimately ruins the estimation accuracy of the swap angle $\theta$ in periodic calibration. In Appendix I, we provide a comprehensive mathematical analysis of periodic calibration and prove that even without complex error and noise, the violation of phase-matching condition makes the estimation variance of $\theta$ scale as $1/\log_{2}(d)$. This is exponentially worse than Heisenberg-limit scaling $1/d^{2}$ when phase-matching condition is satisfied as depth increases. A formal statement can be found in Theorem 23. Moreover, the extremely complex optimization landscape, detailed in Section I.4, renders the estimation using periodic calibration impractical. Though the periodic calibration with phase-matching, i.e. RPE, has higher Fisher information than QSPE, the hardness of satisfying phase-matching condition due to finite resource and time-dependent drift error renders the ideal high accuracy estimation challenging. In contrast, by averaging over $\omega$-points, our QSPE is more robust against error by separating $\theta$ and $\varphi$ estimation processes. This is also empirically justified using real data derived from quantum devices in Figures 4 and 5 in later sections.
@@
    Decoherence. Exploiting the analysis in the Fourier space provides a fruitful structure for mitigating decoherence. To illustrate, we propose a mitigation scheme for the globally depolarizing error in Section G.1. Numerical simulation shows that the scheme can accurately mitigate the depolarizing error and can drastically improve the performance of QSPE estimators.
@@
    Time-dependent noise. We numerically investigate the robustness of the QSPE estimators against realistic qubit frequency-drift error [34] based on observation from real experiments. We show in Section G.3 that the QSPE estimators preserve their accuracy in the presence of this error.
@@
    Readout errors. In Section G.4, we make an explicit resource estimation for a sufficiently accurate mitigation of readout errors.
@@
In this section, we review the experimental deployment of our metrology method and compare it against the leading alternative methods. We consider learning extremely small swap angles in FsimGates, which are important for fermionic simulation and native to Transmon superconducting qubits. FsimGates are two-qubit $U$-gates. The invariant subspace, referred to as single-excitation subspace, is spanned by single-excitation basis |$\ket{01}$ and |$\ket{10}$. CZ gate is a special FsimGate with zero swap angle. Consequently, FsimGates with extremely small swap angles model the imperfect production of CZ gates whose angle parameter estimation is crucial for applications of CZ gate including demonstrating surface code [1]. We refer to Section A.3 for details.
@@
We use the Google Quantum AI superconducting qubits [3] platform to conduct the experiments described in Algorithm D.4 and Fig. 2. We apply our QSPE method to calibrate $\theta$ and $\varphi$ angles of seventeen pairs of CZ gates. Each CZ gate qubit pair is labeled by the coordinates $(x_{1},y_{1})$ and $(x_{2},y_{2})$ of both qubits on a two-dimensional-grid architecture. We plot the statistics of the learned gate angles in Figure 4: the unwanted swap angle for most qubits are small, of order below $10^{-2}$. In comparison, periodic calibration yields unstable estimates with a highly variant standard deviation across different runs (see Figure 7 in Section A.1), a result of its sensitivity to time-dependent errors.
@@
The performance advantage of QSPE over prior art lies in its robustness against time-dependent noise in the single-qubit phase $\varphi$. In traditional methods, such as XEB and robust phase estimation[23], the measurement observables are nonlinear functions of both $\varphi$ and $\theta$, so if there is time-dependent drift in $\varphi$ during each experiment, or over different repetitions of the same experiment routine, the value of inferred $\theta$ will be directly affected (see Figure 7). In comparison, QSPE is tolerant to realistic time-dependent error in $\varphi$ when estimating the swap angle $\theta$ due to the analytic separation between the two parameters through signal transformation, signal processing and Fourier analysis.
@@
To validate the stability of QSPE method, we repeat the same phase estimation routine on each CZ gate pair over 10 independent repetitions. This allows us to bootstrap the variance of the QSPE estimator on $\theta$ and $\varphi$. We show the measured variance and mean of the $\theta$ and $\varphi$ estimates on seventeen CZ-gate pairs over different circuit depths $d$ used in QSPE in Figure 4. We show that on average the variance in $\theta$ estimates is around $10^{-7}$ for a depth-10 QSPE experiment. This corresponds to $3\times 10^{-4}$ in STD, which is one to two orders of magnitude lower than the value of $\theta$ itself. In comparison, we also performed the same set of experiments using XEB and compared the results to QSPE in Figure 5. The variance of $\theta$ inferred by XEB is of order $10^{-4}$ (three orders of magnitude larger than QSPE). Consequently, we show that XEB and periodic calibration are insufficient to learn the value of $\theta$ in our experiments with a larger than unity signal-to-noise ratio.
Figure 4: Learning CZ gate with extremely small unwanted swap angle. Each data point is the average of $10$ independent repetitions and the error bars in the top panels stand for the standard deviation across those repetitions. The number of measurement samples is set to $M=1\times 10^{4}$.
Figure 5: Comparison of the variance in learning swap angle $\theta$ of CZ gates over seventeen pairs of qubits between QSPE and XEB each repeated for 10 times.
@@
In earlier sections, we demonstrate the QSPE algorithm’s effectiveness for estimating angles when the swap angle is of small magnitude. The actual use of QSPE is not limited to this parameter regime. In this subsection, we propose a generalization of QSPE for general swap angles. Theoretical analyses in Appendix D shows that noise in Fourier space is significantly reduced, which consequently suggests the algorithm design using Fourier-transformed data. The key observation is that the exact expression of the amplitude of Fourier coefficients, which is referred to as $A_{k}(\theta)$, can be efficiently solved. Hence, given a set of experimental data, we can estimate the swap angle $\theta$ by aligning experimental amplitudes with theoretical expressions, effectively solving systems of nonlinear equations. When using $d$ $U$-gates per circuit, as discussed in Appendix E, we outline an empirical noise-robust algorithm estimating $\theta$ to error $\epsilon$ using $\mathcal{O}(d\log(d)\epsilon^{-1})$ classical operations. In the numerical results in Appendix E, we demonstrate that the $\theta$-estimation error remains below $5\times 10^{-4}$ for general $\theta$, even with only five $U$-gates. Thanks to Fourier transformation, the angle $\varphi$ is inferred from the phases of Fourier coefficients, decoupling its estimation from $\theta$. It also allows the use of the QSPE $\varphi$-estimator in Equation 9 for varied swap angles, despite the signal-to-noise ratio varies with different swap angles. This is analyzed in the numerical results in Appendix E.
@@
An important of QSPE, thanks to its exceptional sensitivity in measuring extremely small rotations, is in learning quantum crosstalk amplitudes. Quantum crosstalk errors arise from unintended quantum interactions between qubits, which become more problematic when gates operate together. These errors create correlations that are either spatial or temporal, posing a challenge to achieving fault-tolerant quantum computation. In the case of tunable superconducting transmon qubits [24], interactions between two qubits are facilitated by a third “coupler” qubit placed between each pair. This setup allows for the two-qubit interaction to be controlled—turned on or off—by adjusting the coupler qubit’s frequency. Yet, even with control over the coupler qubit’s frequency, there remains a non-zero amount of coupling between neighboring qubits’ different levels. This coupling mimics the system’s Bose-Hubbard coupling Hamiltonian: $H_{\text{crosstalk}}=g_{\text{crosstalk}}\left(\hat{a}_{1}\hat{a}_{2}^{\dagger}+\hat{a}_{1}^{\dagger}\hat{a}_{2}\right)$, where $\hat{a}_{i}$ denotes the bosonic annihilation operator for the $i$-th qubit.
Furthermore, the main effect of quantum crosstalk in qubit subspace can be described by a rotation within the single qubit subspaces $\text{Span}\{\ket{10},\ket{01}\}$ of the two qubits affected by crosstalk. Without any gate operation, the crosstalk’s impact over a period $\Delta t$ follows the same pattern as shown in Eq. (16), where $\theta=g_{\text{crosstalk}}\Delta t$ depends on the crosstalk strength and the duration of crosstalk interaction. This insight allows for the learning of crosstalk effects using QSPE by substituting the $U$-gate in Figure 1 with an idle gate for an appropriate duration $\Delta t$, ensuring that $g_{\text{crosstalk}}\Delta t$ is sufficiently large to be measurable, yet not so large as to compromise the assumptions underlying QSPE. For example, by setting the circuit depth to $d=5$ and $\Delta t$ to 200 ns, the precision in measuring $g_{\text{crosstalk}}$ can reach around 10 mHz, markedly surpassing the accuracy of state-of-the-art results in similar systems, which are around 1 MHz [3].
@@ Discussion
Prior to this work, error mitigation routines had been largely separated from quantum metrology protocols, preventing us from achieving the ultimate limit permitted by physics. Our successful combination of error mitigation with metrology hinges upon treating quantum metrology as a type of quantum signal processing: amplifying a given quantum signal while de-amplifying the unwanted experimental noise. Our work, therefore, opens new directions for using advanced quantum simulation techniques in the design of quantum metrology algorithms in order to achieve novel properties necessary for extremely high accuracy gate learning against realistic environmental noise.
@@
Our future work aims to generalize the optimality of the metrology algorithm against more types of errors and schemes. First, our proof of QSPE estimators’ optimality is based on analyzing the Monte-Carlo sampling error. To fully optimize the design of statistical estimators against all types of error in addition to sampling errors requires modeling and studying the behavior and statistics of all types of dominant realistic error using tools from classical statistics, Bayesian inference, and machine learning. Secondly, here we do not optimize all possible state initialization and measurement schemes. Although theoretical analysis and numerical simulation prove that QSPE estimators are optimal in the given parameter regime and the given state preparation and measurement scheme, it remains an open question whether we can derive the optimal estimators in the most generic setting in 1 by optimizing circuit structure at initialization and measurement steps. Thirdly, the QSPE estimators are designed for the low-depth regime ($d<10$). The restriction to this finite-depth setting is tied in with our main objective of mitigating the detrimental effect of time-dependent noise where deeper circuit depths introduce more drift and decoherence error. This low-depth limit can be lifted if we can fully mitigate various noise effects that kick in at deeper depth. Furthermore, generalizing a deterministic estimator from our work to a variational one can also offer greater flexibility and optimality but requires a deeper understanding of the landscape inherited from the QSPE structure. Lastly, qubitization techniques [18, 10] and cosine-sine decomposition [7, 31] provide powerful two-dimensional subspace representations associated with any unitary matrix. These techniques could potentially generalize our estimation method to large systems with a large number of qubits, which will be our future work.
@@
The experimental data that support the findings are available from the authors upon request.
@@ Acknowledgments
Supplementary Information
Appendix A Preliminaries
A.1 Prior art in quantum gate calibration
@@
A widely used gate calibration technique is called Periodic or Floquet calibration [23, 2], which is an extension of robust phase estimation [14] to multi-parameter regime. It leverages the excitation-preserving structure of the FsimGate to measure the parameters using a restricted set of circuits (compared to full process tomography). This technique amplifies unitary errors in the gate through repeated applications between measurements. When phase-matching condition is attained, it leads to variance in the estimated parameters that scales inversely with the square of the number of gate applications, thus achieving the Heisenberg limit. In [23], two calibration circuit types are discussed, each utilizing different initialization and measurement techniques. They are referred to as the phase method and the population method. The phase method is able to calibrate the phase accumulation angle. However, given the simpler parametric expression, the population method is more frequently used for calibrating swap angles and phase differences [2, 23]. Consequently, in this paper, we focus on the population method for periodic calibration. The quantum circuit is depicted in Figure 6.
@@
Figure 6: Quantum circuit for periodic calibration using the population method.
@@
One difficulty with these techniques is that small values of the swap angle $\theta$ are difficult to be amplified in the presence of larger single-qubit phases. This can be addressed adaptively, by first measuring the unwanted single-qubit phases and applying compensating pulses, but this strategy is limited by the precision with which one can compensate, and the speed with which these single-qubit phases drift relative to the experiment time. For these reasons, in practice estimation of the swap angle is often done with a depth-1 circuit, commonly referred to as unitary tomography [9].
@@
In Appendix I, we provide a comprehensive analysis of periodic calibration. By bounding the Fisher information, we show that the optimal estimation variance depends heavily on the satisfaction of phase-matching condition. Though Heisenberg-limit scaling is achieved with perfect phase-matching, realistic errors, e.g. time-dependent drift errors, render the perfect satisfaction of phase-matching condition challenging. According to the analysis, when the swap angle is small, a slight violation of the phase-matching condition can completely undermine the optimal Heisenberg-limit scaling and renders the estimation variance exponentially worse in terms of depth dependency. These make the use of periodic calibration impractical when the swap angle is small.
@@
Figure 7: Distribution of run-to-run variation of swap-angle estimation across a device. The swap angles were estimated using periodic calibration on four independent datasets for each CZ gate, with 10,000 samples per circuit and a maximum depth of 30. Due to the behavior of the periodic-calibration estimator for particularly small swap angles, a substantial fraction of swap angles were estimated to be identically 0, leading to the portion of the cumulative distribution function that extends off the plot to the left. Discarding these instances leaves us with a median run-to-run standard deviation of close to $4\times 10^{-3}$ radians.
@@
A.2 Quantum signal processing and polynomial analysis
The quantum circuit used in QSPE (see Figure 1) contains a periodic circuit structure in which the $U$-gate and a Z-rotation are interleaved. This circuit structure coincides with a quantum algorithm called quantum signal processing (QSP) [17, 10]. QSP is an useful quantum algorithm for solving numerical linear algebra problems such as quantum linear system problems and Hamiltonian simulation by properly choosing a set of phase factors [8, 22]. Specifically, in this paper, we will use the polynomial structure induced by the theory of QSP [17, 10, 33]. Though the quantum circuit used in the paper is a special case of general QSP circuit by fixing all phase modulation angles, the general QSP structure may provide a more robust paradigm against stochastic phase drift errors by relaxing the fixed angle constraint. Hence, for completeness, we provide a concise overview of the theory of QSP structure in this subsection.
The following theorem is a simplified version of [33, Theorem 1].
@@ A.2 Quantum signal processing and polynomial analysis
Condition (1)”: Note that $\mathrm{SU}(2)$ matrices satisfy
@@
Condition (2)”: To see condition (2), we note that under the transformation $x\mapsto-x$, we have

### v2

@@ Results
The key behind the success of QSPE is the isolation of $\theta$ and $\varphi$ estimations in Fourier space. This enables the robustness of individual angle estimation against the error and noise in another angle. A prior art that is widely used in the gate calibration in Google’s superconducting platform is periodic calibration or Floquet calibration [3, 26]. Periodic calibration measures the transition probability between tensor product states |$\ket{10}$ and |$\ket{01}$ of a periodic quantum circuit with $d$ $U$-gates inside. This differs from our QSPE method which initializes Bell states, though the main body of the quantum circuit is the same. To provide an estimation of angles, periodic calibration uses a black-box optimization to minimize the distance between the parametric ansatz and experimentally measured values. Periodic calibration is based on RPE [16, 26, 3] and generalizes RPE to the estimation of multiple angles. Though RPE provably saturates the Heisenberg limit, the actual performance of periodic calibration highly relies on the satisfaction of the so-called phase-matching condition, namely, $\omega=\varphi$. Previous experiments suggest that the violation of the condition would lower the estimation accuracy of $\theta$ angle by a few magnitudes [26]. Because the phase angle $\varphi$ is vulnerable to time-dependent drift errors, the uncertainty of $\varphi$ ultimately ruins the estimation accuracy of the swap angle $\theta$ in periodic calibration. In Supplementary Note 9, we provide a comprehensive mathematical analysis of periodic calibration and prove that even without complex error and noise, the violation of phase-matching condition makes the estimation variance of $\theta$ scale as $1/\log_{2}(d)$. This is exponentially worse than Heisenberg-limit scaling $1/d^{2}$ when the phase-matching condition is satisfied as depth increases. A formal statement can be found in Theorem 21 in Supplementary Note 9. Moreover, the complex optimization landscape, detailed in Supplementary Note 9 D, renders the estimation using periodic calibration impractical. Though the periodic calibration with phase-matching, i.e. RPE, has higher Fisher information than QSPE, the hardness of satisfying the phase-matching condition due to finite resource and time-dependent drift error renders the ideal high accuracy estimation challenging. In contrast, by averaging over $\omega$ points, our QSPE is more robust against error by separating $\theta$ and $\varphi$ estimation processes. This is also empirically justified using real data derived from quantum devices in Figures 4 and 5 in later sections.
@@
    Decoherence. Exploiting the analysis in the Fourier space provides a fruitful structure for mitigating decoherence. To illustrate, we propose a mitigation scheme for the globally depolarizing error in Supplementary Note 7 A. Numerical simulation shows that the scheme can accurately mitigate the depolarizing error and can drastically improve the performance of QSPE estimators.
@@
    Time-dependent noise. We numerically investigate the robustness of the QSPE estimators against realistic qubit frequency-drift error [38] based on observation from real experiments. We show in Supplementary Note 7 C that the QSPE estimators preserve their accuracy in the presence of this error.
@@
    Readout errors. In Supplementary Note 7 D, we make an explicit resource estimation for sufficiently accurate mitigation of readout errors.
  4. 4.
    Initial state errors. In Supplementary Note 7 E, we analyze the induced estimation error due to the error in the initial states. We demonstrate that these induced estimation errors are negligibly small in real experimental settings.
@@
In this section, we review the experimental deployment of our metrology method and compare it against the leading alternative methods. We consider learning small swap angles in FsimGates, which are important for fermionic simulation and native to Transmon superconducting qubits. FsimGates are two-qubit $U$-gates. The invariant subspace, referred to as single-excitation subspace, is spanned by single-excitation basis |$\ket{01}$ and |$\ket{10}$. CZ gate is a special FsimGate with zero swap angle. Consequently, FsimGates with small swap angles model the imperfect production of CZ gates whose angle parameter estimation is crucial for applications of CZ gate including demonstrating surface code [2]. We refer to Supplementary Note 1 C for details.
@@
We use the Google Quantum AI superconducting qubits [4] platform to conduct the experiments described in Algorithm displayed in Box 1 and Figure 2. We apply our QSPE method to calibrate $\theta$ and $\varphi$ angles of seventeen pairs of qubits on which CZ gates act. Each CZ gate qubit pair is labeled by the coordinates $(x_{1},y_{1})$ and $(x_{2},y_{2})$ of both qubits on a two-dimensional-grid architecture. We plot the statistics of the learned gate angles in Figure 4: the unwanted swap angle for most qubits are small, of order below $10^{-2}$. In comparison, periodic calibration yields unstable estimates with a highly variant standard deviation across different runs (see Supplementary Figure 2 in Supplementary Note 1 A), a result of its sensitivity to time-dependent errors.
@@
The performance advantage of QSPE over prior art lies in its robustness against time-dependent noise in the single-qubit phase $\varphi$. In traditional methods, such as XEB and robust phase estimation [26], the measurement observables are nonlinear functions of both $\varphi$ and $\theta$, so if there is time-dependent drift in $\varphi$ during each experiment, or over different repetitions of the same experiment routine, the value of inferred $\theta$ will be directly affected (see Supplementary Figure 2 in Supplementary Note 1 A). In comparison, QSPE is tolerant to realistic time-dependent error in $\varphi$ when estimating the swap angle $\theta$ due to the analytic separation between the two parameters through signal transformation, signal processing and Fourier analysis.
@@
To validate the stability of QSPE method, we repeat the same phase estimation routine on each CZ gate pair over 10 independent repetitions. This allows us to bootstrap the variance of the QSPE estimator on $\theta$ and $\varphi$. We show the measured variance and mean of the $\theta$ and $\varphi$ estimates on seventeen CZ-gate pairs over different circuit depths $d$ used in QSPE in Figure 4. It is important to note that we apply the technique discussed in Supplementary Note 7 A to mitigate globally depolarizing errors using information from Fourier space. This error mitigation procedure estimates a globally depolarizing circuit fidelity $\alpha$ for each pair of qubits on which CZ gates act, as shown in the right-most column of Figure 4. We observe that the circuit fidelity demonstrates a clear exponential decay with increasing circuit depth, which is consistent with our theoretical analysis [5]. We show that on average the variance in $\theta$ estimates is around $10^{-7}$ for a depth-10 QSPE experiment. This corresponds to $3\times 10^{-4}$ in STD, which is one to two orders of magnitude lower than the value of $\theta$ itself. In comparison, we also performed the same set of experiments using XEB and compared the results to QSPE in Figure 5. The variance of $\theta$ inferred by XEB is of order $10^{-4}$ (three orders of magnitude larger than QSPE). Consequently, we show that XEB and periodic calibration are insufficient to learn the value of $\theta$ in our experiments with a larger than unity signal-to-noise ratio.
@@
In earlier sections, we demonstrate the QSPE algorithm’s effectiveness for estimating angles when the swap angle is of small magnitude. The actual use of QSPE is not limited to this parameter regime. In this subsection, we propose a generalization of QSPE for general swap angles. Theoretical analyses in Supplementary Note 4 show that noise in Fourier space is significantly reduced, which consequently suggests the algorithm design using Fourier-transformed data. The key observation is that the exact expression of the amplitude of Fourier coefficients, which is referred to as $A_{k}(\theta)$, can be efficiently solved. Hence, given a set of experimental data, we can estimate the swap angle $\theta$ by aligning experimental amplitudes with theoretical expressions, effectively solving systems of nonlinear equations. When using $d$ $U$-gates per circuit, as discussed in Supplementary Note 5, we outline an empirical noise-robust algorithm estimating $\theta$ to error $\epsilon$ using $\mathcal{O}(d\log(d)\epsilon^{-1})$ classical operations. In the numerical results in Supplementary Note 5, we demonstrate that the $\theta$-estimation error remains below $5\times 10^{-4}$ for general $\theta$, even with only five $U$-gates. Thanks to Fourier transformation, the angle $\varphi$ is inferred from the phases of Fourier coefficients, decoupling its estimation from $\theta$. It also allows the use of the QSPE $\varphi$-estimator in Equation 9 for varied swap angles, despite the signal-to-noise ratio varies with different swap angles. This is analyzed in the numerical results in Supplementary Note 5.
@@
An important application of QSPE, thanks to its exceptional sensitivity in measuring small rotations, is in learning quantum crosstalk amplitudes. Quantum crosstalk errors arise from unintended quantum interactions between qubits, which become more problematic when gates operate together. These errors create correlations that are either spatial or temporal, posing a challenge to achieving fault-tolerant quantum computation. In the case of tunable superconducting transmon qubits [27], interactions between two qubits are facilitated by a third “coupler” qubit placed between each pair. This setup allows for the two-qubit interaction to be controlled—turned on or off—by adjusting the coupler qubit’s frequency. Yet, even with control over the coupler qubit’s frequency, there remains a non-zero amount of coupling between neighboring qubits’ different levels. This coupling mimics the system’s Bose-Hubbard coupling Hamiltonian: $H_{\text{crosstalk}}=g_{\text{crosstalk}}\left(\hat{a}_{1}\hat{a}_{2}^{\dagger}+\hat{a}_{1}^{\dagger}\hat{a}_{2}\right)$, where $\hat{a}_{i}$ denotes the bosonic annihilation operator for the $i$-th qubit.
Furthermore, the main effect of quantum crosstalk in qubit subspace can be described by a rotation within the single qubit subspaces $\text{span}\{\ket{10},\ket{01}\}$ of the two qubits affected by crosstalk. Without any gate operation, the crosstalk’s impact over a period $\Delta t$ follows the same pattern as shown in Supplementary Equation (4), where $\theta=g_{\text{crosstalk}}\Delta t$ depends on the crosstalk strength and the duration of crosstalk interaction. This insight allows for the learning of crosstalk effects using QSPE by substituting the $U$-gate in Figure 1 with an idle gate for an appropriate duration $\Delta t$, ensuring that $g_{\text{crosstalk}}\Delta t$ is sufficiently large to be measurable, yet not so large as to compromise the assumptions underlying QSPE. For example, by setting the circuit depth to $d=5$ and $\Delta t$ to 200 ns, the precision in measuring $g_{\text{crosstalk}}$ can reach around 10 MHz, markedly surpassing the accuracy of state-of-the-art results in similar systems, which are around 1 MHz [4].
@@ Discussion
Prior to this work, error mitigation routines had been largely separated from quantum metrology protocols, preventing us from achieving the ultimate limit permitted by physics. Our successful combination of error mitigation with metrology hinges upon treating quantum metrology as a type of quantum signal processing: amplifying a given quantum signal while de-amplifying the unwanted experimental noise. Our work, therefore, opens directions for using advanced quantum simulation techniques in the design of quantum metrology algorithms in order to achieve properties necessary for high-accuracy gate learning against realistic environmental noise.
@@
Our future work aims to generalize the optimality of the metrology algorithm against more types of errors and schemes. First, our proof of QSPE estimators’ optimality is based on analyzing the Monte-Carlo sampling error. To fully optimize the design of statistical estimators against all types of error in addition to sampling errors requires modeling and studying the behavior and statistics of all types of dominant realistic error using tools from classical statistics, Bayesian inference, and machine learning. Secondly, here we do not optimize all possible state initialization and measurement schemes. Although theoretical analysis and numerical simulation prove that QSPE estimators are optimal in the given parameter regime and the given state preparation and measurement scheme, it remains an open question whether we can derive the optimal estimators in the most generic setting in 1 by optimizing circuit structure at initialization and measurement steps. Thirdly, the QSPE estimators are designed for the low-depth regime ($d<10$). The restriction to this finite-depth setting is tied in with our main objective of mitigating the detrimental effect of time-dependent noise where deeper circuit depths introduce more drift and decoherence error. A related work employing a dynamical-decoupling-based scheme [13] achieves similar accuracy; however, it is not effective in the low-depth limit of our algorithm and lacks optimality guarantees due to its reliance on decoupling sequences, necessitating deeper circuits. This low-depth limit can be lifted if we can fully mitigate various noise effects that kick in at deeper depth. Furthermore, generalizing a deterministic estimator from our work to a variational one can also offer greater flexibility and optimality but requires a deeper understanding of the landscape inherited from the QSPE structure. Lastly, qubitization techniques [20, 11] and cosine-sine decomposition [8, 34] provide powerful two-dimensional subspace representations associated with any unitary matrix. These techniques could potentially generalize our estimation method to large systems with a large number of qubits, which will be our future work.
Methods
Quantum experiments
Quantum experiments in our work are conducted using the Google Quantum AI superconducting qubits [4] platform. Our QSPE method is detailed in the algorithm shown in Box 1 and Figure 2. The details of the XEB method, used in Figure 5, are discussed in Supplementary Note 1 A.
Numerical simulations
All numerical tests are implemented in python. The numerical studies in Supplementary Note 4 E and Supplementary Note 7 C are conducted using Cirq. The study of our method’s robustness against quantum errors in Supplementary Note 7 C is performed using Cirq’s built-in noisy simulator. Other numerical studies, which do not directly involve quantum circuits, are performed using numpy.
@@ Methods
All data presented in this work are visualized in the figures and tables within the main text and the Supplementary Information file.
@@ Acknowledgments
Algorithm 1  Inferring unknown angles in $U$-gate with small swap angle using QSPE
Initiate a complex-valued data vector $\bm{\mathfrak{h}}\in\mathbb{C}^{2d-1}$.
for $j=0,1,\cdots,2d-2$ do
 Set the tunable $Z$-phase angle as $\omega_{j}=\frac{j}{2d-1}\pi$.
 Set $\mathfrak{h}_{j}\leftarrow p_{X}(\omega_{j})-\frac{1}{2}+\mathrm{i}\left(p_{Y}(\omega_{j})-\frac{1}{2}\right)$.
end for
Compute the Fourier coefficients $\mathbf{c}=\mathsf{FFT}\left(\bm{\mathfrak{h}}\right)$.
Output: Estimators $\hat{\theta},\hat{\varphi}$.
Box 1: Algorithm for inferring unknown angles in $U$-gate with small swap angle using QSPE
Fig. 1: Quantum circuit for QSPE with an exemplified two-qubit $U$-gate. The input quantum state is prepared to be Bell state in either $\ket{+_{\ell}}$ or |$\ket{\mathrm{i}_{\ell}}$ according to the type of experiment. The quantum circuit enjoys a periodic structure of the unknown $U$-gate and a tunable $Z$ rotation.
Fig. 2: Flowchart of main procedures in QSPE. The experimental data are collected from depth $d$ quantum circuit experiments featuring equally-spaced phase modulation angles $\omega$, as shown in the left panels. Probabilities from each experiment of different phase modulations are analyzed using Fourier transformation. As illustrated in the right panels, the Fourier-space data are better structured compared to real-space data. Gate angles are then derived using our QSPE estimators.
Fig. 3: A nontrivial transition of the optimal variance in solving QSPE. The theoretical analysis of the transition is in Supplementary Note 6. a Phase diagram showing the nontrivial transition of the optimal variance in solving QSPE. QSPE estimators attain the optimal variance in the pre-asymptotic regime. b Cramér-Rao lower bound (CRLB) and the theoretically derived estimation variance. The single-qubit phases are set to $\varphi=\pi/16$ and $\chi=5\pi/32$. The number of measurement samples is set to $M=1\times 10^{5}$.
Fig. 4: Learning CZ gate with small unwanted swap angle. Each data point is the average of $10$ independent repetitions and the error bars in the top panels stand for the standard deviation across those repetitions. The number of measurement samples is set to $M=1\times 10^{4}$. These columns display the estimated values of gate angles $\theta,\varphi$, and circuit fidelity $\alpha$.
Fig. 5: Comparison of the variance in learning swap angle $\theta$ of CZ gates over seventeen pairs of qubits between QSPE and XEB each repeated for 10 times.
Supplementary Information for Optimal Low-Depth Quantum Signal-Processing Phase Estimation
Supplementary Information 1 Preliminaries
Supplementary Information 1.1 Prior art in quantum gate calibration
@@
A widely used gate calibration technique is called Periodic or Floquet calibration [26, 3], which is an extension of robust phase estimation [16] to multi-parameter regime. It leverages the excitation-preserving structure of the FsimGate to measure the parameters using a restricted set of circuits (compared to full-process tomography). This technique amplifies unitary errors in the gate through repeated applications between measurements. When phase-matching condition is attained, it leads to variance in the estimated parameters that scales inversely with the square of the number of gate applications, thus achieving the Heisenberg limit. In [26], two calibration circuit types are discussed, each utilizing different initialization and measurement techniques. They are referred to as the phase method and the population method. The phase method is able to calibrate the phase accumulation angle. However, given the simpler parametric expression, the population method is more frequently used for calibrating swap angles and phase differences [3, 26]. Consequently, in this paper, we focus on the population method for periodic calibration. The quantum circuit is depicted in Figure 1.
@@
Supplementary Figure 1: Quantum circuit for periodic calibration using the population method.
@@
One difficulty with these techniques is that small values of the swap angle $\theta$ are difficult to amplify in the presence of larger single-qubit phases. This can be addressed adaptively, by first measuring the unwanted single-qubit phases and applying compensating pulses, but this strategy is limited by the precision with which one can compensate, and the speed with which these single-qubit phases drift relative to the experiment time. For these reasons, in practice estimation of the swap angle is often done with a depth-1 circuit, commonly referred to as unitary tomography [10].
@@
In Section Supplementary Information 9, we provide a comprehensive analysis of periodic calibration. By bounding the Fisher information, we show that the optimal estimation variance depends heavily on the satisfaction of the phase-matching condition. Though Heisenberg-limit scaling is achieved with perfect phase-matching, realistic errors, e.g. time-dependent drift errors, render the perfect satisfaction of phase-matching condition challenging. According to the analysis, when the swap angle is small, a slight violation of the phase-matching condition can completely undermine the optimal Heisenberg-limit scaling and render the estimation variance exponentially worse in terms of depth dependency. These make the use of periodic calibration impractical when the swap angle is small.
@@
Supplementary Figure 2: Distribution of run-to-run variation of swap-angle estimation across a device. The swap angles were estimated using periodic calibration on four independent datasets for each CZ gate, with 10,000 samples per circuit and a maximum depth of 30. Due to the behavior of the periodic-calibration estimator for particularly small swap angles, a substantial fraction of swap angles were estimated to be identically 0, leading to the portion of the cumulative distribution function that extends off the plot to the left. Discarding these instances leaves us with a median run-to-run standard deviation of close to $4\times 10^{-3}$ radians.
@@
Supplementary Information 1.2 Quantum signal processing and polynomial analysis
The quantum circuit used in QSPE (see Figure 1 in the main text) contains a periodic circuit structure in which the $U$-gate and a Z-rotation are interleaved. This circuit structure coincides with a quantum algorithm called quantum signal processing (QSP) [19, 11]. QSP is a useful quantum algorithm for solving numerical linear algebra problems such as quantum linear system problems and Hamiltonian simulation by properly choosing a set of phase factors [9, 25]. Specifically, in this paper, we will use the polynomial structure induced by the theory of QSP [19, 11, 37]. Though the quantum circuit used in the paper is a special case of general QSP circuit by fixing all phase modulation angles, the general QSP structure may provide a more robust paradigm against stochastic phase drift errors by relaxing the fixed angle constraint. Hence, for completeness, we provide a concise overview of the theory of QSP structure in this subsection.
The following theorem is a simplified version of [37, Theorem 1].
@@
“Condition (1)”: Note that $\mathrm{SU}(2)$ matrices satisfy
@@
“Condition (2)”: To see condition (2), we note that under the transformation $x\mapsto-x$, we have